The difference of twice the smaller of two consecutive integers and the larger is 39. find the integers?
Answer by TimothyLamb(4379)

2s - b = 39
b - s = 1
b = 1 + s
---
2s - (1 + s) = 39
s - 1 = 39
Answer:
s = 40
b = 41
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php